What is the oxidation state of nitrogen the following redox reaction?

Cu(s) + HNO3(aq) → Cu2+(aq) + NO(g)
Answer #1


For HNO3:
Oxidation number of O = -2
Oxidation number of H = +1

lets the oxidation number of N be x

use:
1* oxidation number (H) + 3* oxidation number (O) + 1* oxidation number (N) = net charge
1*(+1)+3*(-2)+1* x = 0
-5 + 1 * x = 0
x = 5
So oxidation number of N = +5

Oxidation number of each element in HNO3 is
O=-2
H=+1
N=+5

For NO:
Oxidation number of O = -2

lets the oxidation number of N be x

use:
1* oxidation number (O) + 1* oxidation number (N) = net charge
1*(-2)+1* x = 0
-2 + 1 * x = 0
x = 2
So oxidation number of N = +2

Oxidation number of each element in NO is
O=-2
N=+2


N in HNO3 has oxidation state of +5
N in NO has oxidation state of +2
So, N in HNO3 is reduced to NO
Oxidation state of N changes from +5 to +2